OpenStudy (anonymous):

The relation between degrees Fahrenheit F and degrees Celsius C is described by the function F=9/5C + 32. In the following ordered pairs, the first element is degrees Celsius and the second element is its equivalent in degrees Fahrenheit. Find the unknown measure in each ordered pair. a.(43,m) b. (-12,n) c.(p,12) d.(q,19) HELP PLEASE.

OpenStudy (campbell_st):

ok... so in the ordered pair 43 = C and m = F find F using the formula F = 5/9 C + 32 so substituting will give \[m = \frac{5}{9} \times 43 + 32\] just evaluate to find m. the same process for (b) for (c) and (d) you need to find C given F

OpenStudy (anonymous):

I know how to to get that far but how do you do the 5/9x43?

OpenStudy (campbell_st):

well its simply \[m = \frac{5 \times 43}{9} + 32\] so m = 215/9 + 32

OpenStudy (anonymous):

ohhhhh okay so then m=55.8?

OpenStudy (campbell_st):

thats correct

OpenStudy (anonymous):

thankyou so much!

OpenStudy (anonymous):

& what do you mean by for (c) and (d) you need to find C given F

OpenStudy (campbell_st):

in (c) 12 = 5/9p + 32 you need to solve for p

OpenStudy (anonymous):

yeah so I'd do the same thing you did up top? 12=5p/9+32? or should I subtract 32 on both sides to get -20=5/9p & go on from there?

OpenStudy (campbell_st):

multiply both sides by 9 to get rid of the denominator what do you get

OpenStudy (anonymous):

108=5 x p +32

OpenStudy (anonymous):

then does it become 108= 5p+32?

OpenStudy (anonymous):

then i subtract 32 on both sides

OpenStudy (campbell_st):

I was using you last form of the equation \[-20 = \frac{5p}{9}\] when I said multiply both sides by 9

OpenStudy (anonymous):

ok so i dont do what i just typed to you?

OpenStudy (campbell_st):

ok... you are correct in getting to \[-20 = \frac{5p}{9}\] now multiply both sides by 9

OpenStudy (anonymous):

alright so the answer is -36=p

OpenStudy (campbell_st):

thats correct well done